In a triangle ABC, If s-a/11=s-b/12=s-c/13, thenTan^2 (A/2) is equal to

Grade:11

1 Answers

Saurabh Koranglekar
5 years ago
a/11 = b/12 = c/13

find LCM(a/11,b/12,c/13)

that is how you'll figure sides

and then you can find angle immediately